OpenStudy (anonymous):

Express cos6x in terms of cosx.

OpenStudy (michele_laino):

do you know complex numbers?

OpenStudy (michele_laino):

we can use the formula of \(De\;Moivre\), so we can write this: \[\begin{gathered} {\left( {\cos x + i\sin x} \right)^6} = {\left( {\cos x} \right)^6} + 6i{\left( {\cos x} \right)^5}\left( {\sin x} \right) - 15{\left( {\cos x} \right)^4}{\left( {\sin x} \right)^2} + \hfill \\ \hfill \\ - 20i{\left( {\cos x} \right)^3}{\left( {\sin x} \right)^3} + 15{\left( {\cos x} \right)^2}{\left( {\sin x} \right)^4} + 6i\left( {\cos x} \right){\left( {\sin x} \right)^5} + \hfill \\ \hfill \\ - {\left( {\sin x} \right)^6} \hfill \\ \end{gathered} \]

OpenStudy (jango_in_dtown):

cos 6x= cos (2.3x)=2 (cos 3x )^2-1 =2[4cos^3x - 3cosx ]^2-1 =2[16 cos^6x+9cos^2x-24cos^4x]-1=32cos^6x+18cos^2x-48cos^4x-1

OpenStudy (michele_laino):

next, we have to separate the imaginary part from the real part at the right side

OpenStudy (michele_laino):

I have used this identity: \[{\left( {\cos x + i\sin x} \right)^6} = \cos \left( {6x} \right) + i\sin \left( {6x} \right)\]

OpenStudy (jango_in_dtown):

@vandanaS if you know about complex numbers, you must use @Michele_Laino 's method. Otherwise use mine

OpenStudy (anonymous):

thanks @jango_IN_DTOWN and @Michele_Laino